OpenStudy (juan1857):

Write r^2-2rsintheta=0 in rectangular form.

OpenStudy (kittiwitti1):

I'll give you a hint... y = rsinθ and x = rcosθ

OpenStudy (juan1857):

yeah its r\[r^{2}-2rsin \theta=0\]

OpenStudy (juan1857):

I got up to x^2+y^2-2y=0, but I think its not complete

OpenStudy (juan1857):

ok, can you show me what I should do next

OpenStudy (kittiwitti1):

...I think I'm stuck lol

OpenStudy (juan1857):

XD

OpenStudy (kittiwitti1):

WAIT NO I GOT IT

OpenStudy (kittiwitti1):

Give me a few minutes xD

OpenStudy (juan1857):

ok lol

OpenStudy (kittiwitti1):

okay I think I have it down now\[r^{2}-2r\sin{\theta}=0\]\[r^{2}=2r\sin{\theta}\]As rsinθ = y, \(r^{2}=2y\). so the equation \(r^{2}=x^{2}+y^{2}\) becomes \(2y=x^{2}+y^{2}\)

OpenStudy (kittiwitti1):

Are you following me? xD

OpenStudy (juan1857):

yeah, I see it, do you think that's it?

OpenStudy (kittiwitti1):

Oh. No, the equation \(x^{2}+y^{2}=r^{2}\) only works if your circle's center is the origin. Here the circle's center is (h,k) or \((x-h)^{2}+(y-k)^{2}=r^{2}\) I don't know if it's clear just yet but there's a k value because of the 2y

OpenStudy (juan1857):

ok so what should i do, lol

OpenStudy (kittiwitti1):

So if we move 2y over to the left side we get \(0=x^{2}+y^{2}-2y\). Then you can complete the square for the y-variable

OpenStudy (kittiwitti1):

Do you know how to complete the square?

OpenStudy (juan1857):

I don't think I do, sorry

OpenStudy (kittiwitti1):

It's okay, that's fine. Completing the square goes something like this - if we have equation \(x^{2}+bx=0\) we can add \(c=\left(\frac{b}{2}\right)^{2}\) to both sides to get \(x^{2}+bx+c=c\) (Actually, you're supposed to make the other side 0 but that's another type of problem)

OpenStudy (kittiwitti1):

So here we have \(x^{2}+y^{2}-2y=0\), and you can perform the completing the square action on the y-values

OpenStudy (kittiwitti1):

\[x^{2}+(y^{2}-2y+\left(\frac{2}{2}\right)^{2})=0+\left(\frac{2}{2}\right)^{2}\]

OpenStudy (kittiwitti1):

I think you can do it from here... Just simplify everything ^_^

OpenStudy (juan1857):

ok, Thanks!

OpenStudy (kittiwitti1):

But I'll be here if you need me to explain anything else :D

OpenStudy (kittiwitti1):

No problem (: \(\text{Happy OpenStudying!!}\)

OpenStudy (juan1857):

You Too, XD

OpenStudy (kittiwitti1):

;)